WLCG-OSG-EGEE Ops' Minutes Mon 26 Jan 2009

Summary

Quiet week. A new version of CREAM was included in the gLite 3.1 update 39 released last week.

Attendance

EGEE

  • Asia Pacific ROC: Jason Shih
  • Central Europe ROC: Malgorzata Krakowian
  • OCC / CERN ROC: Steve Traylen, John Shade, Antonio Retico, Diana Bosio, Farida Naz
  • French ROC: Absent
  • German/Swiss ROC: Angela Poschlad
  • Italian ROC: Absent
  • Northern Europe ROC: Ron Trompert
  • Russian ROC: Lev Shamardin, Victor Edneral
  • South East Europe ROC: Kostas Koumantaros
  • South West Europe ROC: Kai Neuffer
  • UK/Ireland ROC: Jeremy Coles
  • GGUS: Torsten Antoni
  • Security: Romain Wartel
  • GGUS Ticket Police: Maria Dimou

WLCG

  • WLCG Service Coordination: Harry Renshall

WLCG Tier 1 Sites

  • ASGC: Jason Shih
  • BNL: Absent
  • CERN site: Absent
  • FNAL: Joe Kaiser, Catalin Dumitrescu
  • FZK: Angela Poschlad
  • IN2P3: Absent
  • INFN: Absent
  • NDGF: Tore Mauset
  • PIC: Kai Neuffer
  • RAL: Derek Ross
  • SARA/NIKHEF: Ron Trompert
  • TRIUMF: Absent

LHC Experiments

  • ATLAS: Alessandro di Girolamo
  • LHCb: Absent
  • CMS: Absent
  • ALICE: Absent

Steve decided that we'd reached quorum at 16:05, so off we went amidst a terrible echo in the conferencing system. [Top-tip: #1 is the magic combination to mute your phone! - JRS]

Feedback on Last Week's Minutes

None was given (but John received a couple of requests for corrections by e-mail, indicating that people do read them!).

EGEE Items

Grid Operator Hand Over on Duty

  Primary Team Secondary Team
From ROC SEE ROC Russia
To ROC CERN ROC DECH

  • Steve checked that the empty reports indeed meant that there was nothing to report. [Although empty reports are fine, it would be better to explicitly state "nothing to report" - JRS]

PPS Reports and Issues

  • Antonio had nothing for this section, so moved straight into...

gLite Release News

  • In a remarkably short intervention, Antonio mentioned that a new version of CREAM was included in the gLite 3.1 update 39 (bug fixes) released last week. Steve asked when other new releases could be expected. Antonio: New VOMS certificates and an FTS patch expected soon (maybe even this week).

EGEE Items From ROC Reports

Please consult the agenda page, since Steve made inline comments which answered most of the issues.
  • On the question of CEs failing SAM tests when the associated close SE was down, Steve confirmed that selecting another close SE was the correct solution, and CE ROC accepted this (they'd wanted to check whether other ROCs had any better ideas).
  • GGUS:45231 has been reassigned to Rutherford, and "they should deal with it".
The terrible echo that was slowly driving Steve mad stopped when, after multiple requests, one of either the anonymous sunrise participant or the Russian ROC muted their phone!
  • Steve requested that SEE elaborate on the BDII stability issue. Kostas replied that he was asking for greater priority be given to this topic.
  • To SWE ROC and the LIP complaint that ATLAS uses 5GB on WNs, Steve pointed out that this is described in the VO cards & that a GGUS ticket should be opened if there's a problem. Xavier Espinal will contact ATLAS. Kai could not find the numbers on the VO-card (Steve remembered that it was something like "space required by Job").

Action on Steve: check VO card setting for this value for all HEP VOs.

  • Steve read out the note from SAM that there's a new version of the portal in Validation which links to GridView for the history display. Please check it out and send any comments to Judit.Novak at CERN.

Operational Security (Romain Wartel)

Romain covered three items:
  • Security incident in Beijing (AP): site responds, but very slowly. CERN needs to help this site to avoid suspension
  • CSIRT mailing list will now be hosted at IN2P3 (rather than CERN). All sites should use new mailing list (due to CERN public mailing lists having problem with SPAM).
  • JSPG recommends that SAM tests should only check that those CAs installed are up-to-date (and not the full list).

WLCG Items

WLCG issues coming from ROC reports

  • None

Upcoming WLCG Service Interventions

  • Consult links on the agenda page. Steve asked whether there were any additional interventions planned for this week. The ensuing silence indicated that there weren't.

WLCG Service Coordination

  • The link to the minutes of the WLCG daily meetings is provided on the agenda page.

No HEP VOs submitted any reports prior to the meeting. Alessandro confirmed that, for ATLAS, nothing of strategic importance had happened.

OSG Items

Before Maria got started with her list, Rob mentioned GGUS:44018 and BNL. A resolved message had been received from BNL, but without any information about how the problem had been resolved. Therefore, it was bounced back to them. Rob explained that OSG don't have the same mechanism for closing tickets as GGUS, but hoped that late spring should see an improvement. Diana nevertheless wondered why the ticket was on hold even though the user had asked that the ticket be closed.

Maria had three tickets on her list:

  • GGUS:44104 : Rob, who was on holiday last week, said there will be an architecture change so ticket stays open for a few more months.
  • GGUS:44140 ATLAS BU problem – no replies from Boston, so Rob will follow up personally
  • GGUS:44837 Rob says that it should be closed

Maria then mentioned GGUS:45094 which is in escalation since today (OSG:6213). The contact at CERN no longer exists. Rob will try and fix this.

Newly Created Action Items

Assigned to Due date Description State Closed Notify  
SteveTraylen 2009-02-02 Check VO-card setting for WN local disk space requirements for all HEP VOs.

Reviewed 3rd February, Alice = 10 GB, Atlas = 15 GB, CMS = 10GB, LHCb = 2 GB.

All LHC VOs specify values for WN disk space. Close this action, if you see particular
VOs exceeding this then submit GGUS tickets for the VO.

Close this after next operations meeting.

2009-02-12 edit

Review of Open Action Items

Open Action Items

IdSubmitterDescriptionCreationDueAssigned To 

Actions Closed in Last 20 Days

IdSubmitterDescriptionCreationDueAssigned ToClosed 

AOB

No one had any.

Next Meeting

The next meeting will be Monday, 02 Feb 2009 15:00 UTC (16:00 Swiss local time).

  • Attendees can join from 14:45 UTC (15:45 Swiss local time) onwards.
  • The meeting will start promptly at 15:00 UTC (16:00 Swiss local time).
  • The WLCG section will start at the fixed time of 15:30 UTC (16:30 Swiss local time).
  • To dial in to the conference:
    • Dial +41227676000
    • Enter access code 0148141


These minutes can only be changed by members of:

Edit | Attach | Watch | Print version | History: r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r4 - 2009-02-12 - SteveTraylen
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    EGEE All webs login

This site is powered by the TWiki collaboration platform Powered by Perl This site is powered by the TWiki collaboration platformCopyright &© by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Ask a support question or Send feedback